OpenStudy (sabrina1):

find point(x,y) on the line y=x using points (-8,10) and(8,-7)

OpenStudy (amistre64):

?

OpenStudy (sabrina1):

ohhh sorry i for got to ADD equidistant point lol

OpenStudy (amistre64):

:) you want the point that is equidistant from the givens and that sits on y=x then right?

OpenStudy (sabrina1):

yes lol :)

OpenStudy (amistre64):

in other words; the place tha thte line cross

OpenStudy (amistre64):

what kind of a line can we create with the 2 given points?

OpenStudy (amistre64):

or am i thinking of this erroneously .....

OpenStudy (amistre64):

we need a distance formula perhaps

OpenStudy (sabrina1):

welll yess i know we need the distance formula however since were lookimg for a set of points what would we replace it by in the equation

OpenStudy (amistre64):

(-8,10) (8,-7) -(x , y) -( x, y) ----------------------- -8-x,10-y 8-x,-7-y these are our respective distances from any point on y=x

OpenStudy (amistre64):

this always gets messy when I try to do it :)

OpenStudy (sabrina1):

LOL :P at least youre honest. i started doing it and threw out half my copybook

OpenStudy (amistre64):

\[\sqrt{x^2+16x+64+y^2-20y+100}\] equals \[\sqrt{x^2-16x+64+y^2+14y+49}\] as long as I did the squareing right that looks to be getting closer

OpenStudy (amistre64):

we can square each one to get rid of the radicals and try to clean it up some: \[x^2+16x+y^2-20y+164=x^2-16x+y^2+14y+113\] \[16x-20y+164=-16x+14y+113\] \[32x-20y+164=14y+113\] \[32x+164=34y+113\] \[32x+51=34y\] hmmm

OpenStudy (sabrina1):

i kinda get it but for some reason im also a little confused lol

OpenStudy (amistre64):

yeah, its trying to use a distance formula to equate the distance from one point to the y=x line; with the other point

OpenStudy (amistre64):

i think where this line and y=x cross is the point of interest tho

OpenStudy (amistre64):

lets try it out :) \[32x+51=34y\] and since y=x \[32x+51=34x\] \[51=2x\] \[51/2=x\] perchance?

OpenStudy (sabrina1):

thats it! thats the answer to it i just wanted to know how to come up with that. such a simple problem turns out to be not so simple. i very much appreciate it:)

OpenStudy (amistre64):

just tough out the distance formulas and equate them \[\sqrt{(x-Px_1)^2+(y-Py_1)^2}=\sqrt{(x-Px_2)^2+(y-Py_2)^2}\]

OpenStudy (sabrina1):

thanks youve been a great help. midterms are the worst! lol

OpenStudy (amistre64):

youre welcome :) good luck